OpenStudy (aashray):

distance between 2 bodies is increased to 3 times of its original value.what is the effect on the gravitational force between them

OpenStudy (anonymous):

You can take ratio of the gravitational force with the 2nd distance to that with the 1st one. In both cases masses m and M are the same, and G is constant: \(\large{\frac{F_{g2}}{F_{g1}} = \frac{G\frac{mM}{(r_2)^2}}{G\frac{mM}{(r_1)^2}}}= G\frac{mM}{(r_2)^2}* \frac{(r_1)^2}{GmM}= \left (\frac{r_1}{r_2}\right)^2\) Since \(r_2\) is 3 times \(r_1\), then \(r_2 = 3r_1\): \(\large{\left (\frac{r_1}{r_2}\right)^2 = \left (\frac{r_1}{3r_1}\right)^2= \left (\frac{1}{3}\right)^2= \frac{1}{9}} \) By increasing the distance 3 times the gravitational force has just \(\frac{1}{9}\) of its original strength.